# MHONGOOSE — A MeerKAT Nearby Galaxy H i Survey

### v1

@@ Acknowledgements.
It is a pleasure to thank SARAO staff members past and present for their help in making MHONGOOSE a reality. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 882793 “MeerGas”). We thank SURF, and particularly Dr. Raymond Oonk for the support in the early stages of the data reduction. KAO is supported by a Royal Society Dorothy Hodgkin Fellowship, by STFC through grant ST/T000244/1, and by the European Research Council (ERC) through Advanced Investigator grant to C.S. Frenk, DMIDAS (GA 786910). DJP, NZ and SK are supported through the South African Research Chairs Initiative of the Department of Science and Technology and National Research Foundation (Grant number 77825). EA and AB thank the CNES for financial support. LC acknowledges funding from the Chilean Agencia Nacional de Investigación y Desarrollo (ANID) through Fondo Nacional de Desarrollo Científico y Tecnológico (FONDECYT) Regular Project 1210992. The work at AIRUB is supported by the German Federal Ministry of Education and Research (BMBF) Verbundforschung grant 05A20PC4 (Verbundprojekt D-MeerKAT-II). KS acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC). LVM and AS acknowledge financial support from the grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and from the grant PID2021-123930OB-C21 funded by MCIN/AEI/10.13039/501100011033, by “ERDF A way of making Europe” and by the European Union. BN acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France.
@@ Appendix A Atlas of single-track moment maps
The second-moment maps are shown using the same intensity scale for all galaxies. For all galaxies, the colour map shows the same velocity range from 0 (light-blue) to 30 (red) km s-1. Contours start at 12 km s-1, and increase in steps of 12 km s-1. For reference, the 24 km s-1 contour is shown in black. Both the first- and second-moment maps have been masked such that only values with ${\rm S/N}>3$ in the zeroth-moment map are shown. While we do not present a full analysis of the first- and second-moment maps in this paper, we do caution that for the edge-on galaxies, these have to be interpreted with care because of the strong projection effects.
@@ Appendix B The effect of beam smearing and inclination on second-moment values
We can compare the distribution shown in Fig. B.18 (right panel) with the much more complex distribution for the real NGC 1566 in Fig. B.18 (left panel). The latter shows higher second-moment values over a much larger range in column densities, indicating that a significant fraction of these higher values are not due to beam smearing. Inspection of Fig. 12 shows that these are rather caused by high-second moment components in galaxies with overall low second-moment values (e.g., J0419–54) as well as galaxies that have overall higher second-moment values (e.g., J1318–21).
We can test this further by taking the inclinations of the galaxies into account. In general, the broadening of profiles due to beam smearing and length of the line of sight through the disk increases dramatically towards high inclinations, with the effect strongest in edge-on galaxies. There is therefore a possibility that high-inclination galaxies contribute to the higher second-moment values we observe at intermediate column densities (i.e., around $\sim 10^{20}$ cm-2). In Fig. B.19 (left panel) we show again the distribution from Fig. 13, which includes all galaxies in the sample, at all inclinations, including edge-ons. To remove the effect that high-inclination galaxies may have on the distribution, we show in the centre panel the distribution limited to galaxies with an inclination $i<60^{\circ}$ (16 galaxies) and in the right panel the distribution limited to inclination $i<70^{\circ}$ (22 galaxies). Comparison of these three distributions shows that omitting the high-inclination galaxies mainly affects the width of the second-moment distribution in the low-column density tail. The effects in the intermediate and high-column density ranges are much less pronounced, and the distribution remains essentially the same in these column density regimes.
Figure A.18: Single-track moment maps for the MHONGOOSE galaxies using the r05_t00 resolution. From left to right: (i): Combined grz-colour image from DECaLS. (ii): primary-beam corrected zeroth-moment or integrated H i intensity map. Contours as indicated in the Figure. The lowest contour represents S/N = 3, with subsequent contour levels increasing by factors of two. (iii): First-moment map or intensity-weighted velocity field. Red colours indicate the receding side, blue colours the approaching side. The central velocity (listed in Tab. 1) is indicated by the thick contour. Other contours are spaced by 10 or 20 km s-1, as indicated in the Figure. (iv): Second-moment map: the colours show the range from 0 (light-blue) to 30 (red) km s-1. The lowest contour shows the 12 km s-1 level, and subsequent contours are spaced by 12 km s-1. The 24 km s-1 contour is shown in black. See text for a more extensive description.
Figure A.18 (Cont.): Single-track moment maps of MHONGOOSE galaxies.
Figure A.18 (Cont.): Single-track moment maps of MHONGOOSE galaxies.
Figure A.18 (Cont.): Single-track moment maps of MHONGOOSE galaxies.
Figure A.18 (Cont.): Single-track moment maps of MHONGOOSE galaxies.

### v2

@@ Acknowledgements.
It is a pleasure to thank SARAO staff past and present for their help in making MHONGOOSE a reality. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation, an agency of the Department of Science and Innovation. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 882793 “MeerGas”). We thank SURF, and particularly Dr. Raymond Oonk for the support in the early stages of the data reduction. KAO is supported by a Royal Society Dorothy Hodgkin Fellowship, by STFC through grant ST/T000244/1, and by the European Research Council (ERC) through Advanced Investigator grant to C.S. Frenk, DMIDAS (GA 786910). DJP, NZ and SK are supported through the South African Research Chairs Initiative of the Department of Science and Technology and National Research Foundation (Grant number 77825). EA and AB thank the CNES for financial support. LC acknowledges funding from the Chilean Agencia Nacional de Investigación y Desarrollo (ANID) through Fondo Nacional de Desarrollo Científico y Tecnológico (FONDECYT) Regular Project 1210992. The work at AIRUB is supported by the German Federal Ministry of Education and Research (BMBF) Verbundforschung grant 05A20PC4 (Verbundprojekt D-MeerKAT-II). KS acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC). LVM and AS acknowledge financial support from the grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 and from the grant PID2021-123930OB-C21 funded by MCIN/AEI/10.13039/501100011033, by “ERDF A way of making Europe” and by the European Union. BN acknowledges financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas A&M University. BASS is a key project of the Telescope Access Program (TAP), which has been funded by the National Astronomical Observatories of China, the Chinese Academy of Sciences (the Strategic Priority Research Program “The Emergence of Cosmological Structures” Grant # XDB09000000), and the Special Fund for Astronomy from the Ministry of Finance. The BASS is also supported by the External Cooperation Program of Chinese Academy of Sciences (Grant # 114A11KYSB20160057), and Chinese National Natural Science Foundation (Grant # 12120101003, # 11433005). The Legacy Survey team makes use of data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), which is a project of the Jet Propulsion Laboratory/California Institute of Technology. NEOWISE is funded by the National Aeronautics and Space Administration. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.
@@ Appendix A Atlas of single-track moment maps
The second-moment maps are shown using the same intensity scale for all galaxies. For all galaxies, the color map shows the same velocity range from 0 (light-blue) to 30 (red) km s-1. Contours start at 12 km s-1, and increase in steps of 12 km s-1. For reference, the 24 km s-1 contour is shown in black. Both the first- and second-moment maps have been masked such that only values with ${\rm S/N}>3$ in the zeroth-moment map are shown. While we do not present a full analysis of the first- and second-moment maps in this paper, we do caution that for the edge-on galaxies, these have to be interpreted with care because of the strong projection effects.
Figure A.18: Single-track moment maps for the MHONGOOSE galaxies using the r05_t00 resolution. From left to right: (i): Combined grz-color image from DECaLS. (ii): primary-beam corrected zeroth-moment or integrated H i intensity map. Contours as indicated in the Figure. The lowest contour represents S/N = 3, with subsequent contour levels increasing by factors of two. (iii): First-moment map or intensity-weighted velocity field. Red colors indicate the receding side, blue colors the approaching side. The central velocity (listed in Tab. 1) is indicated by the thick contour. Other contours are spaced by 10 or 20 km s-1, as indicated in the Figure. (iv): Second-moment map: the colors show the range from 0 (light-blue) to 30 (red) km s-1. The lowest contour shows the 12 km s-1 level, and subsequent contours are spaced by 12 km s-1. The 24 km s-1 contour is shown in black. See text for a more extensive description.
Figure A.18. continued: Single-track moment maps of MHONGOOSE galaxies.
Figure A.18. continued: Single-track moment maps of MHONGOOSE galaxies.
Figure A.18. continued: Single-track moment maps of MHONGOOSE galaxies.
Figure A.18. continued: Single-track moment maps of MHONGOOSE galaxies.
@@ Appendix B The effect of beam smearing and inclination on second-moment values
We can compare the distribution shown in Fig. B.19 (right panel) with the more complex distribution for the real NGC 1566 in Fig. B.19 (left panel). The latter shows higher second-moment values over a larger range in column densities, indicating that a significant fraction of these higher values are not due to beam smearing. Inspection of Fig. 12 shows that these are caused by high-second moment components in galaxies with overall low second-moment values (e.g., J0419–54) as well as galaxies that have overall higher second-moment values (e.g., J1318–21).
We can test this further by taking the inclinations of the galaxies into account. In general, the broadening of profiles due to beam smearing and length of the line of sight through the disk increases dramatically towards high inclinations, with the effect strongest in edge-on galaxies. There is therefore a possibility that high-inclination galaxies contribute to the higher second-moment values we observe at intermediate column densities (i.e., around $\sim 10^{20}$ cm-2). In Fig. B.20 (left panel) we show the distribution from Fig. 13, which includes all galaxies in the sample, at all inclinations, including edge-ons. To remove the effect that high-inclination galaxies may have on the distribution, we show in the centre panel the distribution of galaxies with an inclination $i<60^{\circ}$ (16 galaxies) and in the right panel the distribution with $i<70^{\circ}$ (22 galaxies). Comparison of these distributions shows that omitting the high-inclination galaxies mainly affects the width of the second-moment distribution in the low-column density tail. The effects in the intermediate and high-column density ranges are much less pronounced, and the distribution remains essentially the same in these column density regimes.